OpenStudy (anonymous):

What is the equation of the line in slope-intercept form that passes through the point (–2, 4) and is perpendicular to the line y = –2x + 4?

OpenStudy (aroub):

y=mx+b , m= 1/2 , (-2,-4) -4=1/2(-2)+b -4=-1+b -4+1=b -3=b y=1/2x-3

OpenStudy (anonymous):

y=1/2x+3

OpenStudy (anonymous):

Those Arent In The Options

OpenStudy (aroub):

what options u have ???

OpenStudy (anonymous):

oops. y=1/2x-5

OpenStudy (anonymous):

y = 1/5x + 2 y = -1/5x – 5 y = 1/5x – 2 y = -1/5x + 5

OpenStudy (anonymous):

y=x/2+5